OpenStudy (anonymous):

Medal!!! Solve -3x2 - 4x - 4 = 0.

OpenStudy (anonymous):

I have this so far: a = -3 b = -4 c = -4 x = 4 +- √(-4)^2 – 4 * -3 * -4 / 2(-3) x = 4 +- √-32 / -6 x = 4 +- √-1 * √4 *√8 / -6

OpenStudy (danjs):

you cant fail if you use the quadratic formula thing

OpenStudy (danjs):

i can tell it will not be a real number solution

OpenStudy (anonymous):

I just can't remember what to do next.

OpenStudy (danjs):

\[x = \frac{ 4 \pm \sqrt{16 - 4*(-3)*(-4)} }{ -6 }\]

OpenStudy (danjs):

simplify the root

OpenStudy (danjs):

\[x = \frac{ 4 \pm \sqrt{-32} }{ -6 }\]

OpenStudy (danjs):

recall \[\sqrt{-32} = \sqrt{-1 * 32} = \sqrt{-1}\sqrt{32} = i *\sqrt{16*2} = 4i*\sqrt{2}\]

OpenStudy (anonymous):

Yes, i have gotten that far i have\[x = 4 +- √-1 * √4 *√8 / -6\]

OpenStudy (danjs):

ok, remember i = square root of -1

OpenStudy (danjs):

\[x = \frac{ 4 \pm 4i \sqrt{2} }{ -6 }\]

OpenStudy (danjs):

you can simplify it a bit more, by dividing everything by 2

OpenStudy (danjs):

\[x = \frac{ 2 \pm 2i \sqrt{2} }{ -3 }\]

OpenStudy (danjs):

\[x = \frac{ -2 }{ 3 } \pm \frac{ -2i }{ 3 }\sqrt{2}\]

OpenStudy (danjs):

Do you understand everything?

OpenStudy (anonymous):

I think

OpenStudy (danjs):

You got to this right? \[x = \frac{ 4 \pm \sqrt{-1}\sqrt{4}\sqrt{8} }{ -6 }\]

OpenStudy (danjs):

replace, square root of -1 with imaginary number i. \[\sqrt{8}\sqrt{4} = \sqrt{32} = \sqrt{16}\sqrt{2} = 4\sqrt{2}\] Instead of root 8 times root 4, use root 16 times root 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@DanJS Could you help me with another?

OpenStudy (anonymous):

Solve 5x2 = -30x - 65.

OpenStudy (anonymous):

@DanJS ?

OpenStudy (anonymous):

@wio can you help?